Find the equation of a Parabola with the vertex at (-3,-2) and directrix of X=1
Answer by lwsshak3(11628)

Find the equation of a Parabola with the vertex at (-3,-2) and directrix of X=1
parabola opens leftward:
Its basic equation: (y-k)^2=-4p(x-h), (h,k)=(x,y) coordinates of the vertex
Vertex: (-3,-2)
axis of symmetry: y=-2
p=4 (distance from vertex to directrix on the axis of symmetry)
4p=16
Equation of given parabola:
(y+2)^2=-16(x+3)